### v2

# Deep Learning for Continuous-Time Stochastic Control with Jumps

Abstract
In this paper, we introduce a model-based deep-learning approach to solve finite-horizon continuous-time stochastic control problems with jumps. We iteratively train two neural networks: one to represent the optimal policy and the other to approximate the value function. Leveraging a continuous-time version of the dynamic programming principle, we derive two different training objectives based on the Hamilton–Jacobi–Bellman equation, ensuring that the networks capture the underlying stochastic dynamics. Empirical evaluations on different problems illustrate the accuracy and scalability of our approach, demonstrating its effectiveness in solving complex, high-dimensional stochastic control tasks. Code is available at https://github.com/jdupret97/Deep-Learning-for-CT-Stochastic-Control-with-Jumps.
1 Introduction
A large class of dynamic decision-making problems under uncertainty can be modeled as continuous-time stochastic control problems. In this paper, we introduce two neural network-based numerical algorithms for such problems in high dimensions with finite time horizon and jumps. More precisely, we consider control problems of the form
$\sup_{\alpha}\mathbb{E}\!\left[\int_{0}^{T}f(t,X^{\alpha}_{t},\alpha_{t})dt+F(X^{\alpha}_{T})\right],$ (1)
for a finite horizon $T>0$, where the supremum is over predictable control processes $\alpha=(\alpha_{t})_{0\leq t\leq T}$ taking values in a subset $A\subseteq\mathbb{R}^{m}$. The controlled process $X^{\alpha}$ evolves in a subset $D\subseteq\mathbb{R}^{d}$ according to
$\hskip-2.84526ptdX^{\alpha}_{t}=\beta(t,X^{\alpha}_{t},\alpha_{t})dt+\sigma(t,X^{\alpha}_{t},\alpha_{t})dW_{t}+\int_{E}\hskip-2.84526pt\gamma(t,X^{\alpha}_{t-},z,\alpha_{t})N^{\alpha}(dz,dt),\ \ X^{\alpha}_{0}=x\in D,\hskip-2.84526pt$ (2)
for an initial condition $x\in D$, an $n$-dimensional Brownian motion $W$ and a controlled random measure $N^{\alpha}$ on $E\times\mathbb{R}_{+}$, with $E=\mathbb{R}^{l}\setminus\{0\}$, and suitable functions $\beta\colon[0,T]\times D\times A\to\mathbb{R}^{d}$, $\sigma\colon[0,T]\times D\times A\to\mathbb{R}^{d\times n}$ and $\gamma\colon[0,T]\times D\times E\times A\to\mathbb{R}^{d}$. The functions $f\colon[0,T]\times D\times A\to\mathbb{R}$ and $F\colon D\to\mathbb{R}$ model the running and final rewards, respectively. We assume the controlled random measure is given by $N^{\alpha}(B\times[0,t])=\sum_{j=1}^{M^{\alpha}_{t}}1_{\left\{Z_{j}\in B\right\}}$ for measurable subsets $B\subseteq E$, where $M^{\alpha}$ is a Poisson process with a stochastic intensity of the form $\lambda(t,X^{\alpha}_{t-},\alpha_{t})$ and $Z_{1},Z_{2},\dots$ are i.i.d. $E$-valued random vectors such that, conditionally on $\alpha$, the random elements $W,M^{\alpha}$ and $Z_{1},Z_{2},\dots$ are independent. Our goal is to find an optimal control $\alpha^{*}$ and the corresponding value of problem (1). In view of the Markovian nature of the dynamics (2), we work with feedback controls of the form $\alpha_{t}=\alpha(t,X^{\alpha}_{t-})$ for measurable functions $\alpha\colon[0,T)\times D\to A$ and consider the value function $V:[0,T]\times D\to\mathbb{R}$ given by
$V(t,x)=\sup_{\alpha}\mathbb{E}\!\left[\int_{t}^{T}f(s,X^{\alpha}_{s},\alpha_{s})ds+F(X^{\alpha}_{T})\,\Big|\,X^{\alpha}_{t}=x\right].$ (3)
Under suitable assumptions111see e.g. Soner (1988), $V$ is the unique solution of the following Hamilton–Jacobi–Bellman (HJB) equation
$\partial_{t}V(t,x)+\sup_{a\in A}H(t,x,V,a)=0,\quad V(T,x)=F(x)\,$, (4)
for the Hamiltonian $H:[0,T)\times D\times\mathcal{V}\times A\rightarrow\mathbb{R}$ given [^2] By $\mathcal{V}$ we denote the set of all functions in $C^{1,2}([0,T)\times\mathbb{R}^{d})$ such that the expectation in (5) is finite for all $t,x$ and $a$. by
$H(t,x,V,a)$ $=f(t,x,a)+\beta^{T}(t,x,a)\nabla_{x}V(t,x)+\frac{1}{2}{\rm Tr}\!\left[\sigma\sigma^{T}(t,x,a)\nabla^{2}_{x}V(t,x)\right]$ (5)
$\hskip 62.59605pt+\lambda(t,x,a)\,\mathbb{E}\!\left[V(t,x+\gamma(t,x,Z_{1},a))-V(t,x)\right]\,$,
where $\nabla_{x}V$, $\nabla^{2}_{x}V$ denote the gradient and Hessian of $V$ with respect to $x$. Our approach consists in iteratively training two neural networks to approximate the value function $V$ and an optimal control $\alpha^{*}$ attaining $V$. It has the following features
  • It yields accurate results for high-dimensional continuous-time stochastic control problems in cases where the dynamics of the underlying stochastic processes are known.
  • It can effectively handle a combination of diffusive noise and random jumps with controlled intensities.
  • It can handle general situations where the optimal control is not available in closed form but has to be learned together with the value function.
  • It approximates both the value function and optimal control at any point $(t,x)\in[0,T)\times D$
While methods like finite differences, finite elements and spectral methods work well for solving partial (integro-)differential equations P(I)DEs in low dimensions, they suffer from the curse of dimensionality, and as a consequence, become infeasible in high dimensions. Recently, different deep learning based approaches for solving high-dimensional PDEs have been proposed (Raissi et al., 2017, 2019; Han et al., 2017, 2018; Sirignano & Spiliopoulos, 2018; Berg & Nyström, 2018; Beck et al., 2021; Lu et al., 2021; Bruna et al., 2024). They can directly be used to solve continuous-time stochastic control problems that admit an explicit solution for the optimal control in terms of the value function since in this case, the expression for the optimal control can be plugged into the HJB equation, which then reduces to a parabolic PDE. On the other hand, if the optimal control is not available in closed form, it cannot be plugged into the HJB equation, but instead, has to be approximated numerically while at the same time solving a parabolic PDE. Such implicit optimal control problems can no longer be solved directly with one of the deep learning methods mentioned above but require a specifically designed iterative approximation procedure. For low-dimensional problems with non-explicit optimal controls, a standard approach from the reinforcement learning (RL) literature is to use generalized policy iteration (GPI), a class of iterative algorithms that simultaneously approximate the value function and optimal control (Jacka & Mijatović, 2017; Sutton & Barto, 2018). Particularly popular are actor-critic methods going back to Werbos (1992), which have a separate memory structure to represent the optimal control independently of the value function. However, classical GPI schemes become impractical in high dimensions as the PDE and optimization problem both have to be discretized and solved for every point in a finite grid. This raises the need of meshfree methods for solving implicit continuous-time stochastic control problems in high dimensions with continuous action space. Several local approaches based on a time-discretization of (2) have been explored by e.g. Han & E (2016); Nüsken & Richter (2021); Huré et al. (2021); Bachouch et al. (2022); Ji et al. (2022); Li et al. (2024); Domingo-Enrich et al. (2024a, b). Alternatively, this can also be achieved using popular (deep) RL techniques including policy gradient methods such as A2C/A3C (Mnih et al., 2016), PPO (Schulman et al., 2017) and TRPO (Schulman et al., 2015); Q-learning type algorithms with DQN (Mnih et al., 2013), C51 (Bellemare et al., 2017), see also Wang et al. (2020); Jia & Zhou (2023); Gao et al. (2024); Szpruch et al. (2024); and hybrid approaches with DDPG (Lillicrap et al., 2019) or SAC (Haarnoja et al., 2018). However, these model-free deep RL algorithms do not explicitly take into account the underlying dynamics (2) of the stochastic control problem (1) but instead, solely on sampling from the environment. As a result, they are less accurate in cases where the state dynamics are known (see Figure 3 below). Moreover, this curse of dimensionality inherent in solving high-dimensional PDEs is further exacerbated when dealing with the jump-diffusion (2), as the resulting HJB PIDE (4) requires the numerical evaluation of jump expectations for every space-time point sampled from the domain.
In this paper, we introduce a deep model-based approach for stochastic control problems with jumps which takes the system dynamics (2) into account by leveraging results derived from the HJB equation (4). This removes the need to simulate the underlying jump-diffusion equation and, as a result, avoids discretization errors. Our approach combines GPI with PIDE solving techniques and approximates the value function and optimal control in an actor-critic fashion with two neural networks trained iteratively on sampled data from the space-time domain. It has the advantage that it provides a global approximation of the value function and optimal control available for all space-time points, which can be evaluated rapidly in online applications. We develop two related algorithms. The first one, GPI-PINN, approximates the value function by training a neural network to minimize the residuals of the HJB equation (4), following a physics-inspired neural network (PINN) approach (Raissi et al., 2017, 2019) while leveraging Proposition 3.1 below to avoid the direct computation of the gradient $\nabla_{x}V$ and Hessian $\nabla_{x}^{2}V$ in the Hamiltonian (5). GPI-PINN can thus be viewed as an extension of the method proposed by Duarte et al. (2024), adapted to a finite-horizon setup with time-dependence and a terminal condition in the HJB equation, leading to time-dependent optimal control strategies and value function. It works well in high dimensions for control problems without jumps in the underlying dynamics (2) $(\gamma=0)$, but becomes inefficient in the presence of jumps as it requires computing the jump-expectation $\mathbb{E}\,V(t,x+\gamma(t,x,Z_{1},a))$ at numerous sample points $(t,x)$ in every iteration of the algorithm. To address this, our second algorithm, GPI-CBU, relies on a continuous-time Bellman updating rule to approximate the value function, thereby circumventing the computation of gradients, Hessians and jump-expectations altogether. This makes it highly efficient for high-dimensional stochastic control problems with jumps, even when the control is not available in closed form.
We illustrate the accuracy and scalability of our approach in different numerical examples and provide comparisons with popular RL and deep-learning control methods. Proofs of theoretical results and additional numerical experiments are given in the Appendix.
2 General approach
Let $\alpha\colon[0,T)\times D\to A$ be a feedback control such that equation (2) has a unique solution $X^{\alpha}$ and consider the corresponding value function
$V^{\alpha}(t,x)=\mathbb{E}\!\left[\int_{t}^{T}f(s,X^{\alpha}_{s},\alpha(s,X^{\alpha}_{s-}))ds+F(X^{\alpha}_{T})\,\Big|\,X^{\alpha}_{t}=x\right],\quad(t,x)\in[0,T]\times D.$
Under appropriate assumptions, one obtains the following two results from standard arguments333see e.g. the arguments in the proofs of Theorems 1.3.1 and 2.2.4 in Bouchard (2021)..
Theorem 2.1 (Feynman–Kac Formula).
$V^{\alpha}$ satisfies the PIDE
$\partial_{t}V^{\alpha}(t,x)+H(t,x,V^{\alpha},\alpha(t,x))=0,\quad V^{\alpha}(T,x)=F(x).$
Theorem 2.2 (Verification Theorem).
Let $v\in\mathcal{V}\cap C([0,T]\times D)$ be a solution of the HJB equation (4) such that there exists a measurable mapping $\hat{\alpha}:[0,T)\times D\to A$ satisfying
$\hat{\alpha}(t,x)\in\operatorname*{arg\,max}_{a\in A}H(t,x,v,a)\ \quad\text{for all }(t,x)\in[0,T)\times D$
and the controlled jump-diffusion equation (2) admits a unique solution for each initial condition $x\in D$. Then $v=V$ and $\hat{\alpha}$ is an optimal control.
Based on Theorems 2.1 and 2.2, we iteratively approximate the value function $V$ and optimal control $\alpha^{*}$ with neural networks [^4] Using a $C^{2}$-activation function in the network $V_{\theta}$ ensures that it belongs to $C^{2}([0,T]\times\mathbb{R}^{d})$. $V_{\theta}\colon[0,T]\times D\to\mathbb{R}$ and $\alpha_{\phi}\colon[0,T]\times D\to A$. For given $\alpha_{\phi}$, we train $V_{\theta}$ so as to solve the controlled HJB equation
$\partial_{t}V_{\theta}(t,x)+H(t,x,V_{\theta},\alpha_{\phi}(t,x))=0,\quad V_{\theta}(T,x)=F(x)\,$, (6)
while for given $V_{\theta}$, $\alpha_{\phi}$ is trained with the goal to maximize the Hamiltonian $H(t,x,V_{\theta},\alpha_{\phi}(t,x)).$
In the following, we introduce two different training objectives for updating the value network $V_{\theta}$, leading respectively to the algorithms GPI-PINN and GPI-CBU. GPI-PINN uses a PINN-type loss together with a trick adapted from Duarte et al. (2024) to bypass the explicit computation of gradients and Hessians, whereas GPI-CBU relies on a continuous-time Bellman updating rule with an expectation-free version of the Hamiltonian, thereby also avoiding the computation of the jump-expectations in (5).
3 GPI-PINN
GPI-PINN, described in Algorithm 1 below, relies on a PINN approach to minimize the residuals of the controlled HJB equation (6) in the value function approximation step. To avoid explicit computations of the gradient $\nabla_{x}V_{\theta}(t,x)$ and Hessian $\nabla^{2}_{x}V_{\theta}(t,x)$, which appear in the Hamiltonian (5), we use the following trick, adapted from Duarte et al. (2024).
Proposition 3.1.
Consider a function $v\in\mathcal{V}$ together with a pair $(t,x)\in[0,T)\times D$. Define the function $\psi:\mathbb{R}\to\mathbb{R}$
by
$\psi(h):=\sum_{i=1}^{n}v\hskip-0.85358pt\left(t+\frac{h^{2}}{2n},x+\frac{h}{\sqrt{2}}\sigma_{i}(t,x,a)+\frac{h^{2}}{2n}\beta(t,x,a)\right),\vskip-2.84526pt$
where $\sigma_{i}(t,x,a)$ is the $i^{th}$ column of the $d\times n$ matrix $\sigma(t,x,a)$. Then,
$\psi^{\prime\prime}(0)=\partial_{t}v(t,x)$ $+\beta^{\top}\hskip-1.42262pt(t,x,a)\,\nabla_{x}v(t,x)+\frac{1}{2}\text{Tr}\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,a)\nabla^{2}_{x}v(t,x)\right].$ (7)
Proposition 3.1 makes it possible to replace the computation of gradients and Hessians of $v$ by evaluating the univariate function $\psi^{\prime\prime}(0)$, the cost of which, using automatic differentiation, is a small multiple of $n\cdot cost(v)$.
To formulate GPI-PINN, we need the extended Hamiltonian
$\mathcal{H}(t,x,v,a):=\partial_{t}v(t,x)+H(t,x,v,a),$ (8)
which by Proposition 3.1, can be written as
$\mathcal{H}(t,x,v,a)=\psi^{\prime\prime}(0)+f(t,x,a)+\lambda(t,x,a)\mathbb{E}\big[v(t,x+\gamma(t,x,Z_{1},a))-v(t,x)\big].$
In Algorithm 1, we simplify notations using $\mathcal{H}(t,x,\theta,\phi):=\mathcal{H}(t,x,V_{\theta},\alpha_{\phi}(t,x)).$
Algorithm 1 GPI-PINN
 Initialize admissible weights $\theta^{(0)}$ for $V_{\theta}$ and $\phi^{(0)}$ for $\alpha_{\phi}$. Choose proportionality factors $\xi_{1},\xi_{2}>0$ and set epoch $k=0$.
repeat
  Step 1: Update the value network $V_{\theta^{(k+1)}}$ for a given control $\alpha_{\phi^{(k)}}$ by minimizing the loss
[…]
where
$\hskip-5.69054pt\mathscr{L}_{1}(\theta,\phi)$ $=\xi_{1}\,\mathbb{E}_{(t,x)\sim\mu}\mathcal{H}^{2}(t,x,\theta,\phi)+\xi_{2}\,\mathbb{E}_{x\sim\nu}\left(V_{\theta}(T,x)-F(x)\right)^{2}$ (10)
  Step 2: Update the control network $\alpha_{\phi^{(k+1)}}$ for a given value network $V_{\theta^{(k+1)}}$ by minimizing the loss
[…]
where
$\mathscr{L}_{2}(\theta,\phi)$ $=-\mathbb{E}_{(t,x)\sim\mu}\mathcal{H}\big(t,x,\theta,\phi\big)$ (12)
  $k\leftarrow k+1$
until some convergence criterion is satisfied.
[…]
Since the expectations in $\mathscr{L}_{1}$, $\mathscr{L}_{2}$ and the extended Hamiltonian (8) are typically not available in closed form, we replace them by sample-based estimates. First, we estimate $\mathcal{H}(t,x,\theta,\phi)$ with $\widehat{\mathcal{H}}(t,x,\theta,\phi)$ by approximating the jump-expectation
$\mathbb{E}\,V_{\theta}(t,x+\gamma(t,x,Z_{1},a_{\phi}(t,x)))\approx\frac{1}{J}\sum_{j=1}^{J}V_{\theta}(t,x+\gamma(t,x,z_{j},a_{\phi}(t,x)))$ (13)
for points $(z_{j})_{j=1}^{J}$ in $E$ sampled from the distribution $\mathcal{Z}$ of $Z_{1}$. Then in every gradient step, we approximate $\mathscr{L}_{1}$ and $\mathscr{L}_{2}$ by
$\widehat{\mathscr{L}}_{1}(\theta,\phi^{(k)})=\frac{\xi_{1}}{M_{1}}\sum_{m=1}^{M_{1}}\Big(\widehat{\mathcal{H}}\big(t_{m},x_{m},\theta,\phi^{(k)}\big)\Big)^{2}+\frac{\xi_{2}}{M_{2}}\sum_{m=1}^{M_{2}}\left(V_{\theta}(T,y_{m})-F(y_{m})\right)^{2}$ (14)
and
$\widehat{\mathscr{L}}_{2}(\theta^{(k+1)},\phi)=-\frac{1}{M_{1}}\sum_{m=1}^{M_{1}}\widehat{\mathcal{H}}\hskip-1.42262pt\left(t_{m},x_{m},\theta^{(k+1)},\phi\right),\$(15)
where $(t_{m},x_{m})_{m=1}^{M_{1}}\in[0,T)\times D$ and $(y_{m})_{m=1}^{M_{2}}\in D$ are sampled from $\mu$ and $\nu$, respectively. In every epoch $k=0,\dots,k_{*}-1$, we initialize $\theta^{(k)}_{0}:=\theta^{(k)}$ and make $N_{1}$ gradient steps
$\theta^{(k)}_{i+1}$ $=\theta^{(k)}_{i}-\eta_{1}\nabla_{\theta}\widehat{\mathscr{L}}_{1}(\theta^{(k)}_{i},\phi^{(k)})=\theta^{(k)}_{i}-\frac{2\eta_{1}\xi_{1}}{M_{1}}\sum_{m=1}^{M_{1}}\widehat{\mathcal{H}}\big(t_{m},x_{m},\theta^{(k)}_{i},\phi^{(k)}\big)\nabla_{\theta}\widehat{\mathcal{H}}\big(t_{m},x_{m},\theta^{(k)}_{i},\phi^{(k)}\big)$
$\hskip 126.61476pt-\frac{2\eta_{1}\xi_{2}}{M_{2}}\sum_{m=1}^{M_{2}}\left(V_{\theta^{(k)}_{i}}\hskip-1.13809pt(T,y_{m})-F(y_{m})\right)\nabla_{\theta}V_{\theta^{(k)}_{i}}\hskip-1.13809pt(T,y_{m}),$ (16)
$i=0,\dots,N_{1}-1$, to obtain $\theta^{(k+1)}=\theta^{(k)}_{N_{1}}$. Then, we initialize $\phi^{(k)}_{0}=\phi^{(k)}$ and perform $N_{2}$ gradient steps
$\phi^{(k)}_{i+1}$ $=\phi^{(k)}_{i}-\eta_{2}\nabla_{\phi}\widehat{\mathscr{L}}_{2}(\theta^{(k+1)},\phi^{(k)}_{i})=\phi^{(k)}_{i}+\frac{\eta_{2}}{M_{1}}\sum_{m=1}^{M_{1}}\nabla_{\phi}\widehat{\mathcal{H}}\hskip-1.42262pt\left(t_{m},x_{m},\theta^{(k+1)},\phi^{(k)}_{i}\right)$, (17)
$i=0,\dots,N_{2}-1,$ to get $\phi^{(k+1)}=\phi^{(k)}_{N_{2}}$.
GPI-PINN then yields global approximations $V_{\theta^{(k_{*})}}$ and $\alpha_{\phi^{(k_{*})}}$ of the value function and optimal control on the whole space-time domain $0,T]\times D$. By using Proposition 3.1, it avoids the computation of the gradients and Hessians appearing in the Hamiltonian. However, it still has two drawbacks that make it inefficient for high-dimensional control problems with jumps. First, it has to approximate the jump-expectations $\mathbb{E}\big[V_{\theta}(t_{m},x_{m}+\gamma(t_{m},x_{m},Z_{1},\alpha_{\phi}(t_{m},x_{m})))\big]$ for all sample points $(t_{m},x_{m})$, $m=1,\dots,M_{1}$, in each of the gradient steps (16)–(17) and for every epoch $k=0,1,\dots,k_{*}$, see Eq. (13). Secondly, since the Hamiltonian is already a second-order integro-differential operator, the gradient steps $\nabla_{\theta}\widehat{\mathcal{H}}$ in (16) require the computation of third order derivatives, which is numerically costly.
4 GPI-CBU
GPI-CBU addresses the shortcomings of GPI-PINN by using a value function updating rule based on the expectation-free operator $G_{\zeta}:[0,T]\times D\times E\times\mathcal{V}\times A\to\mathbb{R}$ given by
$G_{\zeta}(t,x,z,v,a)=v(t,x)+\zeta$ $\big[\partial_{t}v(t,x)+f(t,x,a)+\beta^{\top}\hskip-1.42262pt(t,x,a)\,\nabla_{x}v(t,x)$ (18)
$\hskip-79.66771pt+\frac{1}{2}\text{Tr}\hskip-0.85358pt\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,a)\nabla^{2}_{x}v(t,x)\right]+\lambda(t,x,a)\left(v(t,x+\gamma(t,x,z,a))-v(t,x)\right)\big]$
for a scaling factor $\zeta\in\mathbb{R}$.
Proposition 4.1.
Let $X^{\alpha}$ be a solution of the jump-diffusion equation (2) corresponding to a feedback control $\alpha\colon[0,T)\times D\to A$ with associated value function $V^{\alpha}\in C^{1,2}([0,T]\times D)$. For given $t\in[0,T)$, let $Y_{t}$ be a $D$-valued random variable independent of $Z_{1}$ such that $\mathbb{E}\,G_{\zeta}^{2}(t,Y_{t},Z_{1},V^{\alpha},\alpha(t,Y_{t}))<\infty$. Then $V^{\alpha}(t,Y_{t})=g(Y_{t})$ for the Borel measurable function $g:D\to\mathbb{R}$ minimizing the mean squared error
$\mathbb{E}\!\left[\big(g(Y_{t})-G_{\zeta}(t,Y_{t},Z_{1},V^{\alpha},\alpha(t,Y_{t}))\big)^{2}\right].$
Proposition 4.1 suggests to update [^5] By (27), the update rule (19) corresponds to $V_{\theta^{(k+1)}}\hskip-1.42262pt=\hskip-1.42262ptT^{\alpha}V_{\theta^{(k)}}\hskip-3.41432pt:=\hskip-1.99168ptV_{\theta^{(k)}}\hskip-0.85358pt+\hskip-0.85358pt\zeta\mathcal{H}(\cdot,V_{\theta^{(k)}},\alpha(\cdot))$ for the continuous-time Bellman updating (CBU) operator $T^{\alpha}$ associated with the feedback control $\alpha$ and fixed point $T^{\alpha}V^{\alpha}=V^{\alpha}$. the value function parameters according to
$\theta^{(k+1)}=\operatorname*{arg\,min}_{\theta}\mathbb{E}\int_{0}^{T}\Big(V_{\theta}(t,Y_{t})-G_{\zeta}\hskip-0.28453pt\big(t,Y_{t},Z_{1},V_{\theta^{(k)}},\alpha(t,Y_{t})\big)\Big)^{2}dt.$ (19)
By adding a penalty term enforcing the terminal condition, we obtain the recursive scheme
$\theta^{(k+1)}=\operatorname*{arg\,min}_{\theta}\mathscr{L}^{(k)}_{1}(\theta)\vskip-2.84526pt$ (20)
for the loss
$\mathscr{L}^{(k)}_{1}(\theta)\hskip-1.13809pt$ $=\hskip-1.13809pt\xi_{1}\mathbb{E}_{(t,x,z)\sim\mu\otimes\mathcal{Z}}\left(V_{\theta}(t,x)-G_{\zeta}(t,x,z,\theta^{(k)},\phi^{(k)})\right)^{2}+\xi_{2}\mathbb{E}_{x\sim\nu}\left(V_{\theta}(T,x)-F(x)\right)^{2},$
where we use the notation $G_{\zeta}(t,x,z,\theta,\phi):=G_{\zeta}(t,x,z,V_{\theta},\alpha_{\phi}(t,x))$. To implement (20), we approximate $\mathscr{L}^{(k)}_{1}(\theta)$ with
$\hskip-5.69054pt\widehat{\mathscr{L}}^{(k)}_{1}\hskip-1.13809pt(\theta)=\frac{\xi_{1}}{M_{1}}\hskip-0.56905pt\sum_{m=1}^{M_{1}}\hskip-1.42262pt\left(\hskip-0.56905ptV_{\theta}(t_{m},x_{m})\hskip-0.56905pt-\hskip-0.56905ptG_{\zeta}\big(t_{m},x_{m},z_{m},\theta^{(k)},\phi^{(k)}\big)\hskip-0.85358pt\right)^{2}\hskip-0.85358pt+\hskip-1.13809pt\frac{\xi_{2}}{M_{2}}\hskip-1.13809pt\sum_{m=1}^{M_{2}}\hskip-1.70717pt\left(V_{\theta}(T,y_{m})-F(y_{m})\right)^{2}\hskip-2.84526pt\hskip-2.84526pt$ (21)
for $(t_{m},x_{m},z_{m})_{m=1}^{M_{1}}\in[0,T)\times D\times E$ sampled from $\mu\otimes\mathcal{Z}$ and $(y_{m})_{m=1}^{M_{2}}\in D$ sampled from $\nu$. In every epoch $k=0,\ldots,k_{*}-1$, we initialize $\theta^{(k)}_{0}=\theta^{(k)}$ and perform $N_{1}$ gradient steps
$\theta^{(k)}_{i+1}$ $=\hskip-0.7113pt\theta^{(k)}_{i}\hskip-1.42262pt-\eta_{1}\nabla_{\theta}\widehat{\mathscr{L}}^{(k)}_{1}\hskip-0.85358pt(\theta^{(k)}_{i})=\theta^{(k)}_{i}-\frac{2\eta_{1}\xi_{2}}{M_{2}}\sum_{m=1}^{M_{2}}\left(V_{\theta^{(k)}_{i}}\hskip-1.13809pt(T,y_{m})-F(y_{m})\right)\nabla_{\theta}V_{\theta^{(k)}_{i}}\hskip-1.13809pt(T,y_{m})$ (22)
[…]
$i=0,...,N_{1}-1$, to obtain $\theta^{(k+1)}=\theta^{(k)}_{N_{1}}$. To update the control network $\alpha_{\phi}$, we next introduce the operator
$\mathcal{G}(t,x,z,\theta,\phi)$ $:=\partial_{t}V_{\theta}(t,x)+f(t,x,\alpha_{\phi}(t,x))+\beta^{\top}\hskip-1.42262pt(t,x,\alpha_{\phi}(t,x))\,\nabla_{x}V_{\theta}(t,x)$
$\hskip-56.9055pt+\frac{1}{2}\text{Tr}\hskip-0.85358pt\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,\alpha_{\phi}(t,x))\nabla^{2}_{x}V_{\theta}(t,x)\right]+\lambda(t,x,\alpha_{\phi}(t,x))\left(V_{\theta}(t,x+\gamma(t,x,z,\alpha_{\phi}(t,x)))-V_{\theta}(t,x)\right),$
which is an expectation-free version of the extended Hamiltonian $\mathcal{H}$ that, instead of the jump-expectation $\mathbb{E}\,V_{\theta}(t,x+\gamma(t,x,Z_{1},a_{\phi}(t,x)))$, only contains a single jump $V_{\theta}(t,x+\gamma(t,x,z,a_{\phi}(t,x)))$, $z\in E$. GPI-CBU updates the parameters of the control network $\alpha_{\phi}$ according to
$\phi^{(k+1)}=\operatorname*{arg\,min}_{\phi}\mathscr{L}^{(k+1)}_{2}(\phi)$
for the loss
$\mathscr{L}^{(k+1)}_{2}(\phi)=-\mathbb{E}_{(t,x,z)\sim\mu\otimes\mathcal{Z}}\ \mathcal{G}(t,x,z,\theta^{(k+1)},\phi)=-\mathbb{E}_{(t,x)\sim\mu}\mathcal{H}(t,x,\theta^{(k+1)},\phi).\vskip 2.84526pt$
We hence minimize the sample estimate
[…]
by setting $\phi^{(k)}_{0}=\phi^{(k)}$ and making $N_{2}$ gradient steps
$\phi^{(k)}_{i+1}$ $=\phi^{(k)}_{i}-\eta_{2}\nabla_{\phi}\widehat{\mathscr{L}}^{(k+1)}_{2}(\phi^{(k)}_{i})=\phi^{(k)}_{i}+\frac{\eta_{2}}{M_{1}}\sum_{m=1}^{M_{1}}\nabla_{\phi}\mathcal{G}(t_{m},x_{m},z_{m},\theta^{(k+1)},\phi^{(k)}_{i})$, (24)
$i=0,\ldots,N_{2}-1$, to obtain $\phi^{(k+1)}=\phi^{(k)}_{N_{2}}$.
Algorithm 2 GPI-CBU
 Initialize admissible neural weights $\theta^{(0)}$ for $V_{\theta}$ and $\phi^{(0)}$ for $\alpha_{\phi}$. Choose learning rates $\eta_{1},\eta_{2}$, proportionality factors $\xi_{1},\xi_{2}$ and numbers of gradient steps $N_{1},N_{2}$. Set epoch $k=0$.
repeat
  Step 0: Generate $M_{1}$ sample points $(t_{m},x_{m},z_{m})\in[0,T)\times D\times E$ from $\mu\otimes\mathcal{Z}$ and $M_{2}$ sample points $y_{m}\in D$ from $\nu$.
  Step 1: Update $V_{\theta^{(k+1)}}$ by minimizing the loss (21)
[…]
with $N_{1}$ gradient steps (22).
  Step 2: Update $\alpha_{\phi^{(k+1)}}$ by minimizing the loss (23)
[…]
with $N_{2}$ gradient steps (24).
  $k\leftarrow k+1$
until some convergence criterion is satisfied.
return $V_{\theta^{(k)}}$ and $\alpha_{\phi^{(k)}}$ and set $k_{*}\leftarrow k$.
Like GPI-PINN, GPI-CBU leverages Proposition 3.1 to bypass the computation of the gradients $\nabla_{x}V_{\theta}$ and Hessians $\nabla^{2}_{x}V_{\theta}$. In addition, it avoids the costly computation of the jump-expectations $\mathbb{E}\big[V_{\theta}\big(t_{m},x_{m}+\gamma(t_{m},x_{m},Z_{1},\alpha_{\phi}(t_{m},x_{m}))\big)\big]$ at each sample point $(t_{m},x_{m})$, $m=1,\ldots,M_{1}$. Also, when updating the value network using (22), it does not have to compute third-order derivatives like (16) since only $\nabla_{\theta}V_{\theta}$ is needed. This relates to the value update rule (19) of GPI-CBU being recursive as $G_{\zeta}$ now depends explicitly on the value weights $\theta^{(k)}$ computed at the previous epoch, in contrast to the residual-based GPI-PINN. On the other hand, since GPI-PINN averages over different jumps in each update, it exhibits more stable convergence than GPI-CBU; see the numerical results in Section 5 below. The proportionality factors $\xi_{1},\xi_{2}$ and the scaling factor $\zeta$ are hyperparameters. $\xi_{1}$ and $\xi_{2}$ can be fine-tuned following Wang et al. (2022). While Proposition 4.1 holds for any scaling factor $\zeta\in\mathbb{R}$, its choice has an influence on the numerical performance of GPI-CBU. In the numerical experiments of this paper, we set $\zeta=1$ as it provides a good trade-off between convergence speed and accuracy of the improvements in GPI-CBU. On the other hand, negative scaling factors always failed to converge to the true solutions with exploding losses $\widehat{\mathscr{L}}_{1}^{(k)}$ and $\widehat{\mathscr{L}}_{2}^{(k)}$ after only a few epochs. Alternatively, one could consider an adaptive scaling factor $\zeta_{k}>0$ depending on the epoch $k$. More details on hyperparameter fine-tuning are given in Appendix B. Finally, we emphasize that the policy updating rule of GPI-CBU (24) is equivalent to that of GPI-PINN (17), except that it circumvents the computation of the jump-expectations in $\widehat{\cal H}$ as it is now based on the expectation-free operator $\mathcal{G}$.
5 Numerical experiments
In our numerical experiments, we choose the DGM architecture of Sirignano & Spiliopoulos (2018) for both the value and optimal control networks, as it has been shown to empirically improve the PINN performance. Details on the network design and hyperparameters are given in Appendix B.
5.1 Linear-quadratic regulator with jumps
We first consider the linear-quadratic regulator (LQR) problem with jumps
$\inf_{\alpha}\mathbb{E}\left[\int_{0}^{T}c_{1}\,\|\alpha_{s}\|^{2}_{2}\,ds+c_{2}\|X^{\alpha}_{T}\|^{2}_{2}\right],\vskip-0.28453pt$ (25)
where the infimum is over $d$-dimensional predictable
processes $(\alpha_{t})_{0\leq t\leq T}$ and $X^{\alpha}$ is a $d$-dimensional process with controlled dynamics
$dX^{\alpha}_{t}=\alpha_{t}\,dt+\Sigma\,dW_{t}+d\sum_{j=1}^{M^{\alpha}_{t}}Z_{j}\,,\ \ X_{0}=x\in\mathbb{R}^{d}\,$, (26)
for a $d\times d$ matrix $\Sigma$, a $d$-dimensional Brownian motion $W$, a Poisson process $M^{\alpha}$ with intensity $\lambda^{\alpha}_{t}=\Lambda_{1}+\Lambda_{2}\|\alpha_{t}\|^{2}_{2}\,$ for $\Lambda_{1},\Lambda_{2}\geq 0$ and independent i.i.d. mean-zero square integrable $d$-dimensional random vectors $Z_{j}$ with $\mathbb{E}[\|Z_{j}\|^{2}_{2}]=\upsilon$, $j=1,2,\dots$ This problem admits a closed-form solution (see Appendix C), which provides a benchmark for our two numerical algorithms. We report their mean absolute errors with respect to the true solutions $V$ and $\alpha^{*}$, defined as ${\rm MAE}_{V}=\frac{1}{M}\sum_{i=1}^{M}|V_{\theta^{(k_{*})}}\hskip-1.05273pt(t_{i},x_{i})-V(t_{i},x_{i})|$ and ${\rm MAE}_{\alpha}=\frac{1}{M}\sum_{i=1}^{M}\|\alpha_{\phi^{(k_{*})}}\hskip-1.13809pt(t_{i},x_{i})-\alpha^{*}(t_{i},x_{i})\|_{2}$ on a test set of size $M$ uniformly sampled from $0,1]\times[-2.5,2.5]^{d}$.
Figure 2(a) compares ${\rm MAE}_{V}$ and runtimes of GPI-PINN and GPI-CBU as functions of the number of epochs $k$ for 10-dimensional ($d=10$) LQR problems (25) with and without jumps. It can be seen that GPI-PINN and GPI-CBU exhibit similar convergence in the number of epochs. The residual-based approach of GPI-PINN makes it more stable than GPI-CBU; see also Baird (1995). On the other hand, GPI-CBU has a lower computational cost. This is already the case without jumps (left plot of Figure 2(a)) since it avoids the numerical evaluation of third-order derivatives but becomes much more significant with jumps (right plot of Figure 2(a)) as it also circumvents the numerical integration of the jumps.
(a) Comparison of ${\rm MAE}_{V}$ (blue) and runtime in seconds (green) of GPI-PINN (solid line) and GPI-CBU (dashed line) for a 10-dimensional LQR problem without jumps (left) and with jumps (right).
Figure 2 shows results of GPI-CBU applied to a 50-dimensional LQR problem with jumps compared to the analytical solution (30)-(33). While GPI-PINN is inefficient for this problem, GPI-CBU achieves high accuracy in the approximation of the value function as well as the optimal policy. Additional results for up to 150-dimensional LQR problems are reported in Appendix C.
Figure 2: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $t=0$ for $x=(x_{1},0,\dots,0)$ with $x_{1}\in[-2.5,2.5]$ for a 50-dimensional LQR problem with jumps. Orange dotted line: numerical results of GPI-CBU with $\pm 1$ standard deviation given by orange shaded area. Blue line: analytical solution (30)–(33).
Figure 3: $\log{\rm MAE}_{V}$ of different deep-learning methods for a 10-dimensional LQR problem with jumps.
Figure 3 shows the accuracy of GPI-CBU on a 10-dimensional LQR problem with jumps compared with the two popular model-free RL algorithms PPO and SAC as well as the model-based discrete-time approach of Han & E (2016) applied to a time-discretization of the state dynamics (26). It can be seen that in this setup, PPO and SAC cannot compete with the two model-based approaches since they do not explicitly use the dynamics (26) but instead solely rely on sampling from the environment. The method of Han & E (2016) outperforms PPO and SAC but does not achieve the same accuracy as GPI-CBU due to discretization errors and since it does not generalize well to unseen points in the test set. Trying to cover the space-time domain well, we ran the method of Han & E (2016) from several randomly sampled starting points $x_{0}\in D$. But being a local method, it tends to learn the optimal control only along the optimal state trajectories $(t,X^{\alpha^{*}}_{t})_{0\leq t\leq T}$, which results in poor performance in parts of the space-time domain that are not explored well. Additional results are discussed in Appendix C.
5.2 Optimal consumption-investment with jumps
[…]
$\mathbb{E}\!\left[\int_{0}^{T}e^{-\rho s}u(c_{s}Y^{\alpha}_{s})ds+e^{-\rho T}U(Y^{\alpha}_{T})\right]$
for two utility functions [^7] In our numerical experiments, we choose CRRA utility functions. $u,U\colon\mathbb{R}^{+}\to\mathbb{R}$, where $Y^{\alpha}_{t}$ is the wealth process evolving like
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r_{t}+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r_{t})-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\sqrt{\sigma^{i}_{t}}\ \Sigma_{S,i}^{\top}\,dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right]$
for a stochastic interest rate $r_{t}$, expected return rates $\mu^{i}$, stochastic volatilities $\sigma^{i}_{t}$ and stochastic jump intensities $\lambda^{i}_{t}$ whose dynamics are given in Appendix D.2 below. We consider strategies of the form $\pi^{i}(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$ and $c(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$. This problem has $d=2n+2$ state variables. The corresponding HJB equation, given in (42) in Appendix D.2, does not have an analytical solution, but Figure 4 shows that GPI-CBU converges numerically. More details about this consumption-investment problem with stochastic coefficients are provided in Appendix D.2.
In Appendix D.1, we also consider a simplified version of the problem where the coefficients $r_{t},\sigma^{i}_{t},\lambda^{i}_{t}$ are constant. This case can be reduced to a 1-dimensional ODE that can be solved with a standard Runge–Kutta scheme to obtain reference solutions. GPI-CBU generates results that are virtually indistinguishable from the Runge–Kutta results.
Figure 4: Losses $\widehat{\mathscr{L}}^{(k)}_{1}\hskip-1.42262pt(\theta^{(k+1)})$ (left) and $\widehat{\mathscr{L}}^{(k)}_{2}\hskip-1.42262pt(\phi^{(k+1)})$ (right) of GPI-CBU as a function of the epoch $k$. The blue curve in the left plot represents the interior loss of $\widehat{\mathscr{L}}^{(k)}_{1}$ and the orange curve its boundary part, see Eq. (21).
6 Conclusions, limitations and future work
In this paper, we have introduced two iterative deep learning algorithms for solving finite-horizon stochastic control problems with jumps. Both use an actor-critic architecture and train two neural networks to approximate the value function and optimal control, providing global solutions over the entire space-time domain without requiring simulation or discretization of the underlying state dynamics. Our first algorithm, GPI-PINN, works well for high-dimensional problems without jumps but becomes computationally infeasible in the presence of jumps. The second algorithm, GPI-CBU, leverages an efficient expectation-free iteration based on Proposition 4.1 which makes it particularly well-suited for high-dimensional problems with jumps. Both algorithms are model-based. As such, they outperform model-free RL methods in cases where the underlying state dynamics are known. The accuracy and scalability of the two algorithms has been demonstrated in different numerical examples.
A limitation of our approach lies in the need to know the underlying dynamics of the state process, which are not always available in real-world applications. While it is reasonable to assume that physical systems obey known laws of motion, dynamics in economics and finance typically need to be inferred from data. However, in such cases, they can be learned in a preliminary step using e.g. recent model-learning algorithms such as Brunton et al. (2016) or Champion et al. (2019). Once an approximate model has been learned from data, one of the proposed algorithms, GPI-PINN or GPI-CBU, can be applied to efficiently solve the resulting stochastic control problem.
Acknowledgments
Financial support from Swiss National Science Foundation Grant 10003723 is gratefully acknowledged. We thank the reviewers for their helpful comments and constructive suggestions.
Appendix A Proofs
A.1 Proof of Proposition 3.1
Letting $v(\cdot)=v\Big(t+\frac{h^{2}}{2n},x+\frac{h}{\sqrt{2}}\sigma_{i}(t,x,a)+\frac{h^{2}}{2n}\beta(t,x,a)\Big)$, the second-order derivative of $\psi(h):=\sum_{i=1}^{n}v(\cdot)$ is given by
$\psi^{\prime\prime}(h)=$ $\sum_{i=1}^{n}\Bigg(\partial^{2}_{tt}v(\cdot)\frac{h^{2}}{n^{2}}+\frac{h}{n}[\nabla_{tx}v(\cdot)]^{\top}\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)+\frac{1}{n}\partial_{t}v(\cdot)+\frac{1}{n}[\nabla_{x}v(\cdot)]^{\top}\beta(t,x,a)$
$\hskip 25.60747pt+\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)^{\hskip-1.99168pt\top}\hskip-2.84526pt\nabla^{2}_{x}v(\cdot)\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)$
$\hskip 25.60747pt+[\nabla_{tx}v(\cdot)]^{\top}\hskip-2.27621pt\left(\frac{\sigma_{i}(t,x,a)}{\sqrt{2}}+\frac{h}{n}\beta(t,x,a)\right)\frac{h}{n}\Bigg)\,.$
Evaluating it at $h=0$ gives
$\psi^{\prime\prime}(0)=\partial_{t}v(t,x)+\left[\nabla_{x}v(t,x)\right]^{\top}\hskip-2.84526pt\beta(t,x,a)+\frac{1}{2}\text{Tr}\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,a)\nabla^{2}_{x}v(t,x)\right],$
which proves the proposition. ∎
A.2 Proof of Proposition 4.1
Since $Y_{t}$ is independent of $Z_{1}$, one obtains from the definition of $G_{\zeta}$ that
$\mathbb{E}[G_{\zeta}(t,Y_{t},Z_{1},V^{\alpha},\alpha(t,Y_{t}))\mid Y_{t}=x]=V^{\alpha}(t,x)$ $+\zeta\hskip 0.28453pt\Big(\partial_{t}V^{\alpha}(t,x)+f(t,x,\alpha(t,x))\hskip-1.42262pt$
$\hskip-119.50157pt+\beta^{\top}\hskip-1.42262pt(t,x,\alpha(t,x))\,\nabla_{x}V^{\alpha}(t,x)+\hskip-1.42262pt\frac{1}{2}\text{Tr}\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,\alpha(t,x))\nabla^{2}_{x}V^{\alpha}(t,x)\right]$
$\hskip-119.50157pt+\lambda(t,x,\alpha(t,x))\,\mathbb{E}\left[V^{\alpha}(t,x+\gamma(t,x,Z_{1},\alpha(t,x)))-V^{\alpha}(t,x)\right]\Big)$
$\hskip-48.36967pt=V^{\alpha}(t,x)+\zeta\,\mathcal{H}(t,x,V^{\alpha},\alpha(t,x))$ (27)
$\hskip-48.36967pt=V^{\alpha}(t,x),$
where the last equality follows from Theorem 2.1. On the other hand, it is well known that
$\mathbb{E}[G_{\zeta}(t,Y_{t},Z_{1},V^{\alpha},\alpha(t,Y_{t}))\mid Y_{t}]=g(Y)$
for the Borel measurable function $g\colon D\to\mathbb{R}$ minimizing the mean squared error
$\mathbb{E}\!\left[\big(g(Y_{t})-G_{\zeta}(t,Y_{t},Z_{1},V^{\alpha},\alpha(t,Y_{t}))\big)^{2}\right];$
see e.g. Theorem 4.1.15 of Durrett (2019). This concludes the proof. ∎
Appendix B DGM Architecture
Figure 5: DGM architecture for the value neural network with $L=3$ (i.e. 4 hidden layers).
Each DGM layer in Figure 5 for the value network $V_{\theta}$ is as follows
$S^{1}$ $=\sigma_{V}\left(W^{1}\cdot(t,x)+b^{1}\right),$
$Z^{\ell}$ $=\sigma_{V}\left(U^{z,\ell}\cdot(t,x)+W^{z,\ell}\cdot S^{\ell}+b^{z,\ell}\right),\quad\ell=1,\ldots,L,$
$G^{\ell}$ $=\sigma_{V}\left(U^{g,\ell}\cdot(t,x)+W^{g,\ell}\cdot S^{1}+b^{g,\ell}\right),\quad\ell=1,\ldots,L,$
$R^{\ell}$ $=\sigma_{V}\left(U^{r,\ell}\cdot(t,x)+W^{r,\ell}\cdot S^{\ell}+b^{r,\ell}\right),\quad\ell=1,\ldots,L,$
$H^{\ell}$ $=\sigma_{V}\left(U^{h,\ell}\cdot(t,x)+W^{h,\ell}\cdot\left(S^{\ell}\odot R^{\ell}\right)+b^{h,\ell}\right),\quad\ell=1,\ldots,L,$
$S^{\ell+1}$ $=\left(1-G^{\ell}\right)\odot H^{\ell}+Z^{\ell}\odot S^{\ell},\quad\ell=1,\ldots,L,$
$V_{\theta}(t,x)$ $=\sigma_{{V}_{out}}\left(W\cdot S^{L+1}+b\right),$
where the number of hidden layers is $L+1$, and $\odot$ denotes element-wise multiplication. The DGM parameters for the value network are
$\theta=\Big\{\hskip-1.42262ptW^{1},b^{1},\left(U^{z,\ell},W^{z,\ell},b^{z,\ell}\right)_{\ell=1}^{L},$ $\left(U^{g,\ell},W^{g,\ell},b^{g,\ell}\right)_{\ell=1}^{L},$
$\left(U^{r,\ell},W^{r,\ell},b^{r,\ell}\right)_{\ell=1}^{L},\left(U^{h,\ell},W^{h,\ell},b^{h,\ell}\right)_{\ell=1}^{L},W,b\Big\}.$
The number of units in each layer is $N$. $\sigma_{V}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is a twice-differentiable element-wise nonlinearity and $\sigma_{V_{out}}:\mathbb{R}^{N}\to\mathbb{V}$ is the output activation function, also twice-differentiable, where $\mathbb{V}\subseteq\mathbb{R}$ is chosen such as to satisfy possible restrictions on the value function’s output, inferred from the form of the stochastic control problem (e.g. non-negative value function). The control network is designed following the same DGM architecture with $\sigma_{\alpha_{out}}:\mathbb{R}^{N}\rightarrow A$. Throughout the numerical examples of the paper, we consider $L=3$, each one with $N\hskip-0.42677pt=\hskip-0.42677pt50$ neurons, see Figure 5. For the value network, we use $\tanh$ as the activation function $\sigma_{V}$ and softplus for $\sigma_{V_{out}}$. For the control network, we adopt $\tanh$ for $\sigma_{\alpha}$, while the output activation $\sigma_{\alpha_{\text{out}}}$ is linear in Example 5.1 and sigmoid in Example 5.2.
Unless otherwise stated, we take a number of sample points $M_{1},M_{2}$ equal to $256$, a number of gradient steps $N_{1},N_{2}$ equal to $64$, and a maximum number of epochs $k_{*}\leftarrow\min\{k_{*},1500\}$. In practice, we sample a new batch of size $M_{1}+M_{2}$ at each gradient step for Step 1 and 2 of Algorithms 1 and 2, covering a total batch size of $(N_{1}+N_{2})(M_{1}+M_{2})$ per epoch $k$. The network parameters are then updated using the Adaptive Moment Estimation with constant learning rates $\eta_{1},\eta_{2}$ equal to $0.001$. For both GPI-PINN and GPI-CBU, the algorithms are fairly robust to these choices. In our setting, the most critical hyperparameters for convergence are then the proportionality factors $\xi_{1},\xi_{2}$, which we determined using the approach of Wang et al. (2022), together with the scaling rate $\zeta$ (specific to GPI-CBU). In our experiments, we set $\zeta=1$ as it provides a good trade-off between convergence speed and accuracy of the improvements in GPI-CBU. Alternatively, one could consider an adaptive scaling factor $\zeta_{k}$ depending on the epoch $k$, similarly to learning rate schedules. On the other hand, we observe that negative scaling factors always fail to converge to the true solutions with exploding losses $\widehat{\mathscr{L}}_{1}^{(k)}$ and $\widehat{\mathscr{L}}_{2}^{(k)}$ after only a few epochs. Finally, Algorithms 1 and 2 are implemented using TensorFlow and Keras, which are software libraries in Python with reverse mode automatic differentiation. GPU acceleration has been used on a NVIDIA RTX 4090.
Appendix C Linear-quadratic regulator with jumps: detailed results
The value function for the LQR problem with jumps (25) is given by
$V(t,x)=\inf_{\alpha}\mathbb{E}\left[\int_{0}^{T}c_{1}\,\|\alpha_{s}\|^{2}_{2}\,ds+c_{2}\|X^{\alpha}_{T}\|^{2}_{2}\ \Big|\,X^{\alpha}_{t}=x\right].\vskip-0.28453pt$ (28)
The corresponding HJB equation is
$0=\partial_{t}V(t,x)+\frac{1}{2}\text{Tr}\left[\Sigma\Sigma^{\top}\hskip-0.85358pt\nabla^{2}V_{x}(t,x)\right]\hskip-1.13809pt+\inf_{a\in\mathbb{R}^{d}}\Big\{c_{1}\|a\|^{2}_{2}$ (29)
$\hskip 85.35826pt+(\Lambda_{1}+\Lambda_{2}\|a\|^{2}_{2})\,\mathbb{E}\left[V(t,x+Z_{1})-V(t,x)\right]+a^{\top}\hskip-0.56905pt\nabla_{x}V(t,x)\Big\}$
with terminal condition $V(T,x)=c_{2}\|x\|^{2}_{2}$. It is straightforward to show using the verification Theorem 2.2 with HJB equation (29) and the Ansatz
$V(t,x)=\frac{1}{2}h(t)\|x\|^{2}_{2}+g(t)\,$, (30)
that $g$ satisfies
$g(t)=\frac{1}{2}\big(\text{Tr}[\Sigma\Sigma^{\top}]+\Lambda_{1}\hskip 0.28453pt\upsilon)\int_{t}^{T}h(s)ds\,$, (31)
where $h$ solves the non-linear ordinary differential equation
$h^{\prime}(t)=\frac{h^{2}(t)}{2c_{1}+h(t)\hskip 0.56905pt\upsilon\hskip 0.56905pt\Lambda_{2}},\quad h(T)=2c_{2}\,$, (32)
which can be computed efficiently using a numerical method such Runge–Kutta of order $5(4)$ (Dormand & Prince, 1980). The optimal control $\alpha^{*}$ is then given by
$\alpha^{*}(t,x)=-\frac{h(t)x}{2c_{1}+h(t)\hskip 0.56905pt\upsilon\hskip 0.28453pt\Lambda_{2}}\,.$ (33)
In the constant intensity case $\Lambda_{2}=0,\Lambda_{1}>0$, (31) and (32) simplify to
$h(t)=\frac{2c_{1}c_{2}}{c_{1}+c_{2}(T-t)}\quad\text{and}\quad g(t)=(\text{Tr}[\Sigma\Sigma^{T}]+\Lambda_{1}v)\,c_{1}\left(\ln(\frac{c_{2}}{c_{1}}(T-t)+1)\right)\,.$
In this section, we assume for each jump size $Z_{j}\sim\mathcal{N}_{d}(0_{d},\Sigma_{J})$, with $\Sigma_{J}\in\mathbb{R}^{d\times d}$. Hence, we sample according to $\mathcal{Z}$ being a Gaussian distribution the $M_{3}$ points $z_{j}^{(m)}$ of GPI-PINN for each $(t_{m},x_{m})$, $m=1,\ldots,M_{1}$, and the $M_{1}$ points $z_{m}$ of GPI-CBU, see Section 4. Moreover, the frequencies with which various states appear in the sample should be roughly proportional to the probabilities of their occurrence under an optimal policy (Bachouch et al., 2022). As one can guess from (28), it is optimal to drive the process $X^{\alpha}$ towards 0. Therefore, since $\mu(t,x)=\mu(x\mid t)\mu(t)$, we assume $\mu(t)\sim\mathcal{U}_{[0,T]}$ and $\mu(x\mid t)\sim\sqrt{t}\,\mathcal{N}_{d}(0_{d},I_{d})$ to generate the values of $x$ and $t$. Similarly, $\nu(x)\sim\sqrt{T}\,\mathcal{N}_{d}(0_{d},I_{d})$. We then use as parameters of the LQR: $T=1$, $c_{1}=1$, $c_{2}=0.25$ and each element of $\Sigma$, $\Sigma_{J}$ is sampled from $\mathcal{U}_{[0,1]}$.
The three following Figures 7(a)-7-8 are obtained with GPI-CBU in $d=50$ dimensions under the constant intensity case $\Lambda_{1}=0.25$, $\Lambda_{2}=0$.
(a) Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $t=0$ for $x=(x_{1},0,\dots,0)$ with $x_{1}\in[-2.5,2.5]$. Orange dotted line: numerical results of GPI-CBU with $\pm 1$ standard deviation given by orange shaded area ($k_{*}=1500$, $\Lambda_{2}=0$). Blue line: analytical solution (30)-(33).
Figure 7: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $x=\mathbf{1}_{50}$ with $t\in[0,1]$. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1500$, $\Lambda_{2}=0$). Blue line: analytical solution (30)-(33).
Figure 8: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and for $x=(x_{1},\bm{1}_{49})$ with $x_{1}\in[-2.5,2.5]$, obtained from GPI-CBU for the LQR problem with $\Lambda_{2}=0$.
We now consider in the following results the controlled intensity case with $\Lambda_{1}=0,\Lambda_{2}=2$, using the same parameters and sampling procedure as above. Since the intensity now also depends on the control, GPI-PINN becomes even more computationally intensive than in the constant intensity case and of no practical interest in the presence of jumps. Consequently, in the following Figures 910 (and 2), we again focus exclusively on GPI-CBU for a $50$-dimensional LQR problem.
Figure 9: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $x=\mathbf{1}_{50}$ with $t\in[0,1]$. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1500$, $\Lambda_{1}=0$). Blue line: analytical solution (30)–(33).
Figure 10: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and for $x=(x_{1},\bm{1}_{49})$ with $x_{1}\in[-2.5,2.5]$, obtained from GPI-CBU for the LQR problem with $\Lambda_{1}=0$.
These results further validate the accuracy of GPI-CBU, even when considering a controlled intensity for the jump process. As expected, the optimal control $\alpha_{1}^{*}$ is lower in magnitude compared to the constant intensity setting, since it increases the intensity and thus the likelihood of jumps, while the optimal strategy still consists in driving $X^{\alpha}$ towards 0. Again, higher ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics, computed from (30)–(33), tend to occur at the boundaries of the domain, see Figure 10.
Finally, you can find below a table summarizing higher-dimensional results for GPI-CBU in the LQR setting with controlled intensity after $k_{*}=5000$ epochs. This illustrates the scalability of the approach.
Dimensions $d$ $MAE_{V}$ $MAE_{\alpha}$ Loss $\mathscr{L}_{1}$ Loss $\mathscr{L}_{2}$ Time (sec.)
5 0.0023 0.0041 0.0237 -0.1332 6,410
10 0.0025 0.0049 0.0314 -0.1972 8,093
50 0.0147 0.0075 0.1267 -0.4206 16,129
100 0.0492 0.00539 0.6970 -0.4461 24,359
150 0.0979 0.0096 3.7210 -0.4671 33,120
Table 1: GPI-CBU’s performance metrics in function of the state dimension $d$ for the LQR example with controlled intensity after $k_{*}=5000$ epochs.
Appendix D Optimal consumption-investment problem
D.1 Constant coefficients
We first study the optimal consumption-investment problem with constant coefficients, where we assume the risk-free asset evolves according to
$dS^{0}_{t}=rS^{0}_{t}dt\,$, (34)
for a constant interest rate $r\in\mathbb{R}$, and with $n$ stocks whose price follows for $i=1,\ldots,n$,
$\frac{dS^{i}_{t}}{S^{i}_{t-}}=\mu^{i}dt+\Sigma^{\top}_{i}dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)$ (35)
for some constants $\mu^{i}\in\mathbb{R}$, $\Sigma_{i}\in\mathbb{R}^{n}$, correlated [^8] Note that $\Sigma=(\Sigma_{1},\ldots,\Sigma_{n})^{\top}\hskip-2.27621pt\in\mathbb{R}^{n\times n}$ can be seen as the upper Choleski decomposition of the correlation matrix $\Sigma\Sigma^{\top}$ of the Brownian motion $W$. $n$-dimensional Brownian motions $W$, independent Poisson processes $M^{i}$ with constant intensities $\lambda^{i}\geq 0$ and i.i.d. normal random vectors $Z^{i}_{1},Z^{i}_{2},...,Z^{i}_{M^{i}_{t}}\sim\mathcal{N}(\mu^{i}_{Z},\sigma^{i}_{Z})$ with $\mu^{i}_{Z}\in\mathbb{R},\sigma^{i}_{Z}\in\mathbb{R}^{+}$. Suppose an investor starts with an initial endowment of $Y_{0}>0$, consumes at rate $c_{t}Y^{\alpha}_{t-}$ and invests in the $n$ stocks according to $\pi^{i}_{t}Y^{\alpha}_{t-}$, $i=1,\dots n$. If the risk-free asset is used to balance the transactions, the resulting wealth evolves according to
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r)-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\Sigma^{\top}_{i}dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right].$ (36)
Let us assume she attempts to optimize
$\mathbb{E}\!\left[\int_{0}^{T}e^{-\rho s}u(c_{s}Y^{\alpha}_{s})ds+U(Y^{\alpha}_{T})\right]\,$,
for two utility functions $u,U\colon\mathbb{R}^{+}\to\mathbb{R}$ and $\rho\in\mathbb{R}$. Since the driving noise in (36) has stationary independent increments, it is enough to consider strategies of the form $c_{t}=c(t,Y^{\alpha}_{t-})$ and $\pi^{i}_{t}=\pi^{i}(t,Y^{\alpha}_{t-})$ for functions $c,\pi^{i}\colon[0,T)\times\mathbb{R}^{+}\to(0,1)$, $i=1,\dots,n$. This $n+1$-dimensional control is then written $\alpha_{t}=(\pi^{1}_{t},\ldots,\pi^{n}_{t},c_{t})^{\top}\hskip-2.84526pt\in A:=(0,1)^{n+1}$. Assuming constant relative risk aversion (CRRA) utility functions $u,U$ with $\gamma\in(0,1)$, the reward functional is then denoted
$V^{\alpha}(t,y)=\mathbb{E}\left[\int_{t}^{T}e^{-\rho(s-t)}\frac{\left(c_{s}Y^{\alpha}_{s}\right)^{\gamma}}{\gamma}ds+\frac{\left(Y^{\alpha}_{T}\right)^{\gamma}}{\gamma}\,\Big|\,Y^{\alpha}_{t}=y\right].$ (37)
Finally, writing $\mu=(\mu^{1},\ldots,\mu^{n})^{\top}$, $\mu_{Z}=(\mu_{Z}^{1},\ldots,\mu_{Z}^{n})^{\top}$, $\sigma_{Z}=(\sigma^{1}_{Z},\ldots,\sigma^{n}_{Z})^{\top}$, $\lambda=(\lambda^{1},\ldots,\lambda^{n})^{\top}$, $\pi=(\pi^{1},\ldots,\pi^{n})^{\top}\hskip-1.42262pt$, the associated HJB equation for the value function $V(t,y)=\sup_{\alpha}V^{\alpha}(t,y)$ satisfies [^9] Note that the term $\rho V(t,y)$ in (38) is not contained in our general HJB equation (4). But GPI-PINN and GPI-CBU still work if it is added. for all $(t,y)\in[0,T)\times\mathbb{R}^{+}$,
$0=$ $\partial_{t}V(t,y)-\rho V(t,y)+\sup_{(c,\pi)\in A}\bigg\{(r+(\mu-r)^{\top}\pi-c)\,y\,\partial_{y}V(t,y)$ (38)
$\hskip-6.25958pt+\frac{1}{2}\pi^{\top}\hskip-1.42262pt\Sigma\Sigma^{\top}\hskip-1.42262pt\pi\,y^{2}\,\partial^{2}_{yy}V(t,y)+\sum_{i=1}^{n}\lambda^{i}\mathbb{E}\left[V(t,y+y\,\pi^{i}(e^{Z^{i}_{1}}-1))-V(t,y)\right]+\frac{(cy)^{\gamma}}{\gamma}\bigg\}\,$,
with $V(T,y)=y^{\gamma}/\gamma$. This stochastic control problem does not admit an analytical solution but we can instead characterize its solution in terms of a PIDE, so as to assess the accuracy of the proposed Algorithms 1 and 2. Following Øksendal & Sulem (2007), we assume the following form $V(t,y)=A(t)y^{\gamma}/\gamma$ for the value function. First order condition on the HJB equation (38) then implies that optimal fractions of wealth is a vector of constants $\pi^{*}(t,y)=\pi^{*}:=(\pi^{1,*},\ldots,\pi^{n,*})^{\top}$ satisfying
$(\mu-r)^{\top}+\Sigma\Sigma^{\top}\pi^{*}(\gamma-1)+\sum_{i=1}^{n}\lambda^{i}\,\mathbb{E}\left[\left(1+\pi^{i,*}(e^{Z^{i}_{1}}-1)\right)^{\gamma-1}\left(e^{Z^{i}_{1}}-1\right)\right]=0\,$, (39)
and the optimal consumption rate $c^{*}$ is independent from the wealth $y$ and given by
$c^{*}(t,y)=A(t)^{1/(\gamma-1)}\,.$ (40)
Plugging back these optimal controls inside the HJB equation (38), we find that the function $A(t)$ satisfies the ODE
$\frac{\partial_{t}A(t)}{A(t)}=\rho\hskip 1.42262pt$- $\gamma\left(r+(\mu-r)^{\top}\pi^{*}-A(t)^{1/(\gamma-1)}\right)-\frac{1}{2}\gamma(\gamma-1)\pi^{*\top}\Sigma\Sigma^{\top}\pi^{*}$ (41)
$\hskip 59.75078pt-\sum_{i=1}^{n}\lambda^{i}\mathbb{E}\left[\left(1+\pi^{i,*}(e^{Z^{i}_{1}}-1)\right)^{\gamma}\hskip-1.42262pt-1\right]-A(t)^{1/(\gamma-1)},$
with $A(T)=1$, which can be solved numerically using again the Runge–Kutta method of order $5(4)$. For both Algorithms 1 and 2, we sample the time and space points independently from the uniform distributions $\mathcal{U}_{[0,T]}$ and $\mathcal{U}_{[0,y_{b}]}$, respectively. The parameters of the optimal investment problem below are as follows: $T=1,y_{b}=150,r=0.02,\rho=0.045,\gamma=0.3,\lambda=0.45\cdot\mathbf{1}_{n},\mu_{Z}=0.25\cdot\mathbf{1}_{n},\sigma_{Z}=0.2\cdot\mathbf{1}_{n},\mu=0.032\cdot\mathbf{1}_{n},\Sigma\Sigma^{\top}=0.2\cdot\mathbf{1}_{n\times n}$ with $\text{diag}(\Sigma\Sigma^{\top})=\mathbf{1}_{n}$. Note that, compared to the LQR problem in Section 5.1, the proportionality factor $\xi$ is now set to $10$. Moreover, since $A=(0,1)^{n+1}$, $\sigma_{\alpha_{out}}$ is chosen to be the sigmoid activation function. Instead of the DGM architecture, we here train a classical feedforward neural network with 4 hidden layers, each of 128 neurons. Finally, the ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics are again defined as in Section 5.1 on a test set of size $M$, uniformly sampled from $0,1]\times[0,y_{b}]$ with $V$ and $\alpha^{*}$ obtained from the RK45 method described above.
We again compare in Figure 11 the ${\rm MAE}_{V}$ and runtime between GPI-PINN and GPI-CBU in function of the number of epochs $k$ for $n=10$ stocks in the consumption-investment problem (37) with and without jumps. We again see that the residual-based GPI-PINN Algorithm 1 tends to be more stable and accurate for larger epochs, despite having a higher runtime. When accounting for jumps in the dynamics (36), GPI-PINN becomes numerically very time-consuming, even for a 10-dimensional control problem. This issue is even amplified compared with the LQR problem (see Figure 2(a)), as the jump size now depends on the control $\pi$ in the wealth dynamics (36). In contrast, GPI-CBU again handles these jumps far more efficiently.
Figure 11: Comparison of the ${\rm MAE}_{V}$ (blue) and the runtime in seconds (green) between GPI-PINN (solid line) and GPI-CBU (dashed line) in function of the number of epochs $k$ for the optimal consumption-investment problem without jumps (left) and with jumps (right) for $n=10$ stocks.
We then address a $50$-dimensional optimal consumption-investment problem with jumps, where only GPI-CBU is implemented as GPI-PINN becomes numerically infeasible. Figures 12 and 13 again confirm the accuracy of GPI-CBU for both the value function and the optimal consumption rate $c^{*}(t)$, with the standard deviation across 10 independent runs of GPI-CBU being virtually imperceptible. The optimal wealth allocations $\pi^{*}$, being constant, are not depicted. However, the ${\rm MAE}_{\alpha}$ heatmap in Figure 14 corroborates our method’s accuracy in determining both $c^{*}$ and $\pi^{*}$. We also observe in this Figure 14 that the higher absolute errors tend to occur at the boundaries of the domain.
Figure 12: Value function $V(t,y)$ at $t=0$ for $y\in[0,150]$ (left) and at $y=50$ for $t\in[0,1]$ (right), with $n=50$ stocks. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$). Blue line: numerical solution from the Runge-Kutta scheme applied to (40)-(41).
Figure 13: Optimal consumption rate $c^{*}(t)$ for $t\in[0,1]$. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$). Blue line: numerical solution from the Runge-Kutta scheme applied to (40)-(41).
Figure 14: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and for $y\in[0,150]$, obtained from GPI-CBU for the optimal consumption-investment problem with constant coefficents.
D.2 Stochastic coefficients
We then study a more complex optimal consumption-investment problem in a realistic market of the form (34)–(35) with a stochastic interest rate
$dr_{t}=a_{r}(b_{t}-r_{t})dt+\nu_{r}\sqrt{r_{t}}\,\Sigma_{r}^{\top}dW_{t},$
with stochastic (Heston) volatility models for the $n$ stock prices
$\frac{dS^{i}_{t}}{S^{i}_{t-}}$ $=\mu^{i}dt+\sqrt{\sigma^{i}_{t}}\,\Sigma^{\top}_{S,i}\,dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\,$,
$d\sigma^{i}_{t}$ $=a^{i}_{\sigma}(b^{i}_{\sigma}-\sigma^{i}_{t})+\nu^{i}_{\sigma}\sqrt{\sigma^{i}_{t}}\,\Sigma^{\top}_{\sigma,i}\,dW_{t},$
and with stochastic jump intensities
$d\lambda^{i}_{t}=a^{i}_{\lambda}(b^{i}_{\lambda}-\lambda^{i}_{t})+\nu^{i}_{\lambda}\sqrt{\lambda^{i}_{t}}\,\Sigma^{\top}_{\lambda,i}\,dW_{t}\,.$
Note that $\Sigma=(\Sigma_{S,1},\ldots,\Sigma_{S,n},\Sigma_{\sigma,1},\ldots,\Sigma_{\sigma,n},\Sigma_{\lambda,1},\ldots,\Sigma_{\lambda,n},\Sigma_{r})^{\top}\in\mathbb{R}^{(3n+1)\times(3n+1)}$ is the (upper) Choleski decomposition of the correlation matrix $\Sigma\Sigma^{\top}$ associated with the $(3n+1)$-dimensional Brownian motions $W$. In this case, the wealth process can still be described as follows
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r_{t}+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r_{t})-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\sqrt{\sigma^{i}_{t}}\ \Sigma_{S,i}^{\top}\,dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right],$
and the strategies should be of the form $c(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$ and $\pi^{i}(t,\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$, with $\alpha_{t}=(\pi^{1}_{t},\ldots,\pi^{n}_{t},c_{t})^{\top}\hskip-2.84526pt\in A:=(0,1)^{n+1}$. Hence, denoting the $(2n+2)$-dimensional process $X^{\alpha}_{t}=(\sigma_{t},\lambda_{t},r_{t},Y^{\alpha}_{t-})$, its dynamics are given by
$dX^{\alpha}_{t}=\mu_{X}(t,X^{\alpha}_{t})dt+\Sigma_{X}(t,X^{\alpha}_{t})dW_{t}+\gamma_{X}(t,X^{\alpha}_{t})\,d\sum_{i=1}^{n}\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right),$
where
$\mu_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}a_{\sigma}(b_{\sigma}-\sigma_{t})\\ a_{\lambda}(b_{\lambda}-\lambda_{t})\\ a_{r}(b_{r}-r_{t})\\ Y^{\alpha}_{t-}(r_{t}+(\mu-r_{t})^{\top}\pi_{t}-c_{t})\end{array}\right),\,\Sigma_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}\nu_{\sigma}\,\sqrt{\sigma_{t}}\,\Sigma^{\top}_{\sigma}\\ \nu_{\lambda}\,\sqrt{\lambda_{t}}\,\Sigma^{\top}_{\lambda}\\ \nu_{r}\,\sqrt{r_{t}}\,\Sigma^{\top}_{r}\\ Y^{\alpha}_{t-}\sum_{i=1}^{n}\pi^{i}_{t}\sqrt{\sigma^{i}_{t}}\,\Sigma_{S,i}^{\top}\end{array}\right),\,$
and
$\gamma_{X}(t,X^{\alpha}_{t})=\left(\begin{array}[]{c}\bm{0}_{n}\\ \bm{0}_{n}\\ 0\\ Y^{\alpha}_{t-}\pi_{t}^{\top}\mathbf{1}_{n}\end{array}\right),$
with $\mu_{X}(\cdot)\in\mathbb{R}^{2n+2},\Sigma_{X}(\cdot)\in\mathbb{R}^{(2n+2)\times(3n+1)}_{+}$ and $\gamma_{X}(\cdot)\in\mathbb{R}^{2n+2}$. Therefore, the value function
$V(t,x)=\sup_{\alpha}\mathbb{E}\left[\int_{t}^{T}e^{-\rho(s-t)}\frac{\left(c_{s}Y^{\alpha}_{s}\right)^{\gamma}}{\gamma}ds+\frac{\left(Y^{\alpha}_{T}\right)^{\gamma}}{\gamma}\,\Big|\,X^{\alpha}_{t}=x\right],$
satisfies the following HJB equation for all $(t,x)\in[0,T)\times\mathbb{R}^{n}_{+}\times\mathbb{R}^{n}_{+}\times\mathbb{R}\times\mathbb{R}^{+}$,
$0=\hskip 2.13394pt$ $\partial_{t}V(t,x)-\rho V(t,x)+\sup_{(c,\pi)\in A}\bigg\{\mu_{X}^{\top}\hskip-0.56905pt(t,x)\,\nabla_{x}V(t,x)+\frac{1}{2}\text{Tr}\left[\Sigma_{X}(t,x)\Sigma_{X}^{\top}(t,x)\,\nabla^{2}_{x}V(t,x)\right]$
$\hskip 18.77882pt+\sum_{i=1}^{n}\lambda^{i}\,\mathbb{E}\left[V(t,\sigma,\lambda,r,y+y\,\pi^{i}(e^{Z^{i}_{1}}-1))-V(t,x)\right]+\frac{(cy)^{\gamma}}{\gamma}\bigg\}\,$, (42)
with terminal condition $V(T,x)=y^{\gamma}/\gamma$. We consider a portfolio of $n=25$ stocks, resulting in a $52$-dimensional value function and a $26$-dimensional control process, with the same parameters as in Section D.1. This version of the consumption-investment problem is significantly more complex than the one discussed in Section D.1, with the value function’s dimensionality increasing from 2 to 52. Consequently, the RK45 numerical method can no longer serve as a reference point to solve the associated HJB equation (42), making it impossible to compute the ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics. However, despite this lack of a reference for comparison, GPI-CBU still produces results that appear reasonable. We indeed first observe in Figure 4 (main manuscript) that the losses $\widehat{\mathscr{L}}^{(k)}_{1}$ (21) and $\widehat{\mathscr{L}}^{(k)}_{2}$ (23) converge as the number of epoch $k$ increases. Note that the orange curve in the left plot represents the interior loss (first right-hand term) of $\widehat{\mathscr{L}}^{(k)}_{1}$, while the blue curve is the boundary loss (second right-hand term) of $\widehat{\mathscr{L}}^{(k)}_{1}$, see Eq. (21).
The following Figures 15 and 16 confirm that the results from GPI-CBU for both the value function and the optimal control are reasonable, as they closely resemble those from Figures 12 and 13. For the value functions of Figure 15, the standard deviations across 10 independent runs of GPI-CBU remain very low, confirming the stability of the approximations. For the optimal consumption rate in Figure 16 (left plot), the standard deviation across the 10 runs tends to be higher for values of time $t$ close to 0, although being still reasonable. Finally, varying the first dimension of the intensity $\lambda^{1}$ mainly impacts the corresponding fraction of wealth $\pi^{1,*}_{t},$ while its effect on the other proportions and consumption rate is more moderate (since arising from the correlation matrix $\Sigma\Sigma^{T}$ of the Brownian motion $W$).
Figure 15: Value function $V(t,x)$ at $t=0$ for $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,y)$ with $y\in[0,150]$ (left) and at $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,50)$ for $t\in[0,1]$ (right). Blue line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$).
Figure 16: Left plot: optimal consumption rate $c^{*}(t,x)$ at $x=(0.15\cdot\mathbf{1}_{10},\mathbf{1}_{10},0.02,50)$ for $t\in[0,1]$ with numerical results from GPI-CBU in blue line and $\pm$ one standard deviation in orange shaded area ($k_{*}=1000$). Right plot: optimal consumption rate $c^{*}(t,x)$ and first five optimal fractions of wealth $\pi^{*}(t,x)$ at $t=0$ and $x=(0.15\cdot\mathbf{1}_{10},\lambda^{1},\mathbf{1}_{9},0.02,50)$ for $\lambda^{1}\in[0,5]$.